OpenStudy (anonymous):

Find a quadratic equation with roots (4+i) and (4-i).

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Whenever you need to find an equation with roots r and s, simply do this: \[\Large f(x) = (x-r)(x-s)\] where r = 4+i and s = 4-i

OpenStudy (anonymous):

okay I'm stuck at the point where I have to multiply... [x-(4+i)][x-(4-i)].... I don't know how to use the foil method for that .

terenzreignz (terenzreignz):

Just foil (x-r) and (x-s) first, and the replace afterwards.

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

okay I get \[x^2-8x+16-i^2\]

terenzreignz (terenzreignz):

Yup... you're almost done ^_^

OpenStudy (anonymous):

\[-i^2 \] ^^^ is equal to 1...?

terenzreignz (terenzreignz):

That's correct :D

OpenStudy (anonymous):

so then the equation is \[x^2 -8x+17=0\]

terenzreignz (terenzreignz):

Bravo ^_^

OpenStudy (anonymous):

okay thank you! it was really that foiling part that had me stuck, cause I knew the method for finding equations given the roots. Didn't realize I could just replace the (x-r)(x-s) with the correct terms... so thanks for pointing that out!

terenzreignz (terenzreignz):

No problem. :)

OpenStudy (anonymous):

ax^2 + bx + c = 0 with 2 complex roots: (4 + i) and (4 - i) Product of 2 roots x1*x2 = (4 + i)(4 - i) = 16 + 1 = 17 = ac Suppose a = 1 and c = 17. Sum of 2 roots: (4 + i) + (4 -i) = 8 = -b. Then b = -8. Back to the equation: x^2 - 8x + 17 = 0 D = b^2 - 4ac = -4 = 4i^2. The 2 roots are:(4 + i) and (4 -i). Correct.